OVHCloud Network Status

Current status
Legend
  • Operational
  • Degraded performance
  • Partial Outage
  • Major Outage
  • Under maintenance
Fibre BHS <> Vaudreuil
Incident Report for Network & Infrastructure
Resolved
Nous avons une coupure fibre confirmée entre le DC de BHS et le site de Vaudreuil. Ceci impact nos circuits optiques vers Newark qui sont passés automatiquement en protection mais aussi les circuits BHS <> Ashburn, BHS <> Toronto et BHS <> Chicago.

We have got a fiber cut confirmed between the Beauharnois datacenter and Vaudreuil. This impact our optical circuits towards Ashburn, Toronto and Chicago

Update(s):

Date: 2018-09-06 12:13:23 UTC
All circuits are now up.

We are waiting the confirmation that all works are finished

Date: 2018-09-06 05:42:32 UTC
The installation of the second section continues with several difficulty.

Modification of the ETR, a gradual restoration of services is now expected from 5:00am local time


Date: 2018-09-06 01:59:48 UTC
Repair works are still ongoing. Initial impact seems to have been caused by vandalism. A new cable is being installed on 2 sections. Cable installation has been completed on first section. Second section is currently in progress.


A new ETR has been provided : fiber pairs continuity will be restored from 12:30 AM local time

Date: 2018-09-05 15:00:21 UTC
The fibercut is localised on a bridge at Valleyfield, QC.
There are about 400 fibers down.
The provider needs to rolled out 3km of new cable to repair.
They just had authorizations from MTQ to begin the work on the bridge.

ETR is 9:00 PM Local Time.


Date: 2018-09-05 05:20:12 UTC
According to the latest update received from our on-site teams. They just finished the trace tests on the first point. They go to the second point to complete the tests to accurately determine the breakpoint. We will keep you informed of developments as soon as we receive new information.
Posted Sep 05, 2018 - 01:30 UTC
This incident affected: Datacenters || BHS (BHS1, BHS2, BHS3, BHS4, BHS5, BHS6, BHS7).